If you buy 5 pitas and 3 salads you get .10 change back from $20. If you buy 8 pitas and 2 salads, it will cost you $26.10. How much change would you get from $20.00 if you bought 3 pitas and 5 salads? (show all steps) Remember to us "let" statement to begin.
Answer by rfer(12670)

let p = pitas
let s = salads
5p+3s=19.90
8p+2s=26.10
-------------
5p+3(26.10-8p)/2=19.90
5p=(78.30-24p)/2=19.90
5p+39.15-12p=19.90
-7p=-19.75
p=2.75
---------------------
5(2.75)+3s=19.90
13.75+3s=19.90
3s=6.15
s=2.05
-----------------------
3p+5s=
3(2.75)+5(2.05)=
8.25+10.25=18.50
20.00-18.50=$1.50 change